Calculate ATP production for the following fatty acids a. heneicosylic acid (21:0) b. lauric acid...

calculate ATP production for the following fatty acids
a. heneicosylic acid (21:0)
b. lauric acid (12:0)
c. dihomo-y-linolenic acid (20:3 Delta 8,11,14)
d. nervonic acid (24:1 Delta 15)
Answer #1

Formulae:

n= no. of carbon atoms in a fatty acid

No. of acetyl CoA formed (N)= \frac{n}{2}

No. of rounds of \beta oxidation of fatty acid = N-1

Each round of \beta oxidation yields=

  • 1 FADH2 = 2 ATP
  • 1 NADH = 3 ATP
  • 1 Acetyl CoA= 12 ATP
  • Total = 17 ATP/round

a. heneicosylic acid (21:0)

This is an odd carbon chained fatty acid with 21 carbons and no double bonds. Here number of acetyl CoA formed will be 9. With this, in addition 1 Succinyl CoA will also be formed which is a 3 Carboned compound.

Thus number of ATP formed by 9 rounds = 9 x 17 ATP/round = 153 ATP

   For 1 Succinyl CoA = 5 ATP

therefore, total ATP = 153+ 5 = 158 ATP

Now, an expenditure of 2 ATP molecules is required to put CoA on the acyl group hence total of 2 ATP will be substracted from yield

Hence total ATP formed = 158-2 = 156 ATP

b. Lauric acid (12:0)

Lauric acid is an even carboned fatty acid with 12 carbons and no double bonds.

Here, No. of acetyl CoA formed will be = 12/2 = 6

No. of rounds of oxidation = 6-1 = 5

No. of ATP= 5 rounds x 17 ATP/round

= 85 ATP molecules + 12 ATP(from 1 acetyl CoA from last step)

   = 97 ATP

Total ATP obtained = 97-2 = 95 ATP

c. dihomo-y-linolenic acid (20:3 Delta 8,11,14)

Here, The fatty acid is 20 carboned with 3 unsaturations.

Thus no. of Acetyl CoA = 20/2 = 10

no. rounds = 10-1 = 9

Total ATP = 9x 17 ATP/rounds = 153 ATP +12 ATP (from remaining 1 Acetyl CoA)

   = 165 ATP

Now, one unsaturation will lead to loss of 2 ATP (3 unsaturations will result in loss of 6 ATP) and 2 ATPs will be required to put CoA on acyl group, hence total of 8 ATPs will be substracted from yield

Total ATP= 165-8 = 157 ATP

d.nervonic acid (24:1 Delta 15)

Here, The fatty acid is 24 carboned with one unsaturation.

Thus no. of Acetyl CoA = 24/2 = 12

no. rounds = 12-1 = 11

Total ATP = 11x 17 ATP/rounds = 187 ATP +12 ATP (from remaining 1 Acetyl CoA)

   = 199 ATP

Now, one unsaturation will lead to loss of 2 ATP and 2 ATPs will be required to put CoA on acyl group, hence total of 4 ATPs will be substracted from yield

Total ATP= 199-4 = 195 ATP